An employer gives a pre-employment evaluation to a large group of applicants.
The scores for the evaluation are normally distributed with a mean of 154 and a standard deviation of 21.


a) What percentage of applicants will score more than 160 on the evaluation?


b) The employer wants to interview only those applicants who score in the top 15%.
What should the cut-off score be for the interviews? Round to the nearest whole number.
Answer #1

Question- L = 21 X~ N(154, 21) 154) N10,1) (a) – fx (x >160) = 13(x=154 > 160-154) - Be (2>0:29) = 0.5- Px COCZC0:29)= 0.5 (b